OpenStudy (anonymous):

what is the slope of a line that passes trough the points( -2 5)and (1 4)

jimthompson5910 (jim_thompson5910):

use the slope formula \[\Large m = \frac{y_{2}-y_{1}}{x_{2}-x_{1}}\]

jimthompson5910 (jim_thompson5910):

(x1,y1) = (-2,5) is your first point (x2,y2) = (1,4) is your second point

OpenStudy (anonymous):

-1/3?

jimthompson5910 (jim_thompson5910):

yes you got it

OpenStudy (anonymous):

thank you im glad im getting better at this

OpenStudy (anonymous):

could you help with another

jimthompson5910 (jim_thompson5910):

me too and sure, go ahead

OpenStudy (anonymous):

the pair of points (6 y and 10 -1 lie on a line with slope 1/4 what is the value of y

jimthompson5910 (jim_thompson5910):

you mean (6,y) and (10,-1) right?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

so you would use the slope formula to get this \[\Large m = \frac{y_{2}-y_{1}}{x_{2}-x_{1}}\] \[\Large \frac{1}{4} = \frac{-1-y}{10-6}\] solve for y

jimthompson5910 (jim_thompson5910):

in this case, (x1,y1) = (6,y) and (x2,y2) = (10,-1)

OpenStudy (anonymous):

but theres not another y

jimthompson5910 (jim_thompson5910):

what do you mean?

OpenStudy (anonymous):

its just 6 y not and another number

jimthompson5910 (jim_thompson5910):

y is a placeholder for some number

jimthompson5910 (jim_thompson5910):

for now, we just leave it as y

jimthompson5910 (jim_thompson5910):

that is why I have -1-y in the numerator in my steps above

OpenStudy (anonymous):

okay im slowly following

OpenStudy (anonymous):

so i would do the same seps just without another number so i would use y as 0

jimthompson5910 (jim_thompson5910):

I'm not sure what you mean

OpenStudy (anonymous):

y is nothing so i just solve it as if y was 0?

jimthompson5910 (jim_thompson5910):

you need to isolate y in this equation \[\Large \Large \frac{1}{4} = \frac{-1-y}{10-6}\]

OpenStudy (anonymous):

how do i do that

jimthompson5910 (jim_thompson5910):

cross multiply, then isolate y

OpenStudy (anonymous):

10-6 =4 -1-y=?

jimthompson5910 (jim_thompson5910):

kinda, but you're missing something

OpenStudy (anonymous):

what is that wait dont i have to divide?

jimthompson5910 (jim_thompson5910):

at some step, I think you do

jimthompson5910 (jim_thompson5910):

but not right away

OpenStudy (anonymous):

so what comes before division

jimthompson5910 (jim_thompson5910):

what do you get when you cross multiply

OpenStudy (anonymous):

ummm 10 y and positive 6 im sorry if i keep getting this wrong

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

it might help to simplify 10-6 first

OpenStudy (anonymous):

-2

jimthompson5910 (jim_thompson5910):

no idk how you're getting that

OpenStudy (anonymous):

just forget it ill medal and fan anyway sorry i don't understand thank you awhole great deal for trying to help me though i really appreciate it

jimthompson5910 (jim_thompson5910):

what is the -2 referring to? I asked you to simplify 10-6 and you got -2 so idk how you got that